# OpsEval: A Comprehensive IT Operations Benchmark Suite for Large Language Models

### v3

@@
Abstract.
Information Technology (IT) Operations (Ops), particularly Artificial Intelligence for IT Operations (AIOps), is the guarantee for maintaining the orderly and stable operation of existing information systems. According to Gartner’s prediction, the use of AI technology for automated IT operations has become a new trend. Large language models (LLMs) that have exhibited remarkable capabilities in NLP-related tasks, are showing great potential in the field of AIOps, such as in aspects of root cause analysis of failures, generation of operations and maintenance scripts, and summarizing of alert information. Nevertheless, the performance of current LLMs in Ops tasks is yet to be determined. A comprehensive benchmark is required to optimize LLMs tailored for Ops (OpsLLM). In this paper, we present OpsEval, a comprehensive task-oriented Ops benchmark designed for LLMs. For the first time, OpsEval assesses LLMs’ proficiency in various crucial scenarios at different ability levels. The benchmark includes 7184 multi-choice questions and 1736 question-answering (QA) formats in English and Chinese. By conducting a comprehensive performance evaluation of the current leading large language models, we show how various LLM techniques can affect the performance of Ops, including self-consistency, chain-of-thought, and few-shot in-context learning, and discussed findings related to various topics, including model quantification, QA evaluation, and hallucination issues. To ensure the credibility of our evaluation, we invite dozens of domain experts to manually review our questions. At the same time, we have open-sourced 20% of the test QA to assist current researchers in preliminary evaluations of their OpsLLM models. The remaining 80% of the data, which is not disclosed, is used to eliminate the issue of the test set leakage. Additionally, we have constructed an online leaderboard that is updated in real-time and will continue to be updated, ensuring that any newly emerging LLMs will be evaluated promptly. Both our dataset and leaderboard have been made public [^1] Data page is available at https://github.com/NetManAIOps/OpsEval-Datasets.
    Leaderboard is available at https://opseval.cstcloud.cn/content/leaderboard.
copyright: noneconference: ; booktitle: ;conference: ; ; ccs: Networks Network managementccs: Computing methodologies Artificial intelligence
1. Introduction
@@
In recent years, large language models (LLMs) have witnessed significant advancements. The latest models, such as GPT-4V (OpenAI, 2023b), GPT-4 (OpenAI, 2023a), LLaMA-2 (et.al., 2023), and ChatGLM3 (202, 2024b), have demonstrated exceptional generalization and task planning capabilities. As a result, these models have provided numerous opportunities to enhance downstream domain-specific applications. With its advanced summarizing, report analyzing, and ability to make API calls, LLM is well suited for Ops on tasks like answer generation, information summarizing, and report analysis. Hereinafter, we refer to the LLM used for Ops as OpsLLM, regardless of whether they have been optimized specifically for Ops.
@@
While there are benchmarks for assessing general-purpose NLP-related capabilities, such as C-EVAL (Huang et al., 2023) and AGIEval (Zhong et al., 2023), there are also benchmarks for specific domains, like FinEval (Zhang et al., 2023) in the financial sector and CMB (Wang et al., 2023a) in the medical field. However, no benchmark exists that can evaluate the effectiveness of LLMs or OpsLLMs in Ops tasks. There is an urgent need for an Ops benchmark that informs us about the performance of current LLMs on Ops tasks. On the other hand, a good benchmark can significantly aid the optimization process of OpsLLMs tailored for the Ops domain.
Nevertheless, due to the specialty of the Ops tasks, constructing an Ops benchmark presents the following challenges:
@@
  • The Ops data is sensitive primarily and proprietary to companies, with very little publicly available data, making it difficult for any company to independently provide a sufficient amount of evaluation data to ensure confidence in the test results.
  • There are many different subdomains (such as 5G communications, cloud computing, log analysis, and network management) and sub-tasks (such as network script generation and terminology explanation) within the Ops field. Different domains and tasks have different accuracy requirements for models. For example, network script generation requires high accuracy, whereas operational terminology explanation may not be as precise as the network script generation task. Automatically annotating evaluation problems that have different accuracy requirements poses a significant challenge.
@@
  • Since existing LLMs are not explicitly trained for the Ops domain, the evaluation results are more sensitive to prompt engineering. Designing appropriate prompts for robust and accurate evaluation is challenging.
@@
  • Existing metrics like BLEU only consider the similarity of model output to ground-truth answers in natural language, which does not always indicate a good response in Ops tasks. Some terms and expressions in Ops scenarios have specific meanings that cannot be summarized in advance. Designing an automatic evaluation method that assesses the accuracy of Q&A from an accurate semantic level is challenging.
@@
To address these issues, we propose OpsEval, a comprehensive IT operations benchmark for LLMs. First, to tackle the challenge of benchmark data being mostly private and not publicly shareable, spread across various companies, we built a community around AIOps with currently 10 companies participating, allowing each entity in the community to continuously contribute a small amount of data. We then aggregate data under the same domain to ensure robustness in evaluation. Additionally, based on publicly available network management books, we generated both multi-choice and QA questions as supplements. To address the challenge of different accuracy requirements across multiple Ops scenarios and tasks, we employed model-based pre-clustering and manual review to annotate eight tasks and three abilities for independent evaluation. In response to the sensitivity of benchmark results to prompts, we systematically test model performance under self-consistency, chain-of-thought, and few-shot in-context learning. The prompts used in our evaluation are also disclosed in the paper. Lastly, to address the inaccuracy of existing metrics like BLEU in assessing the performance of question answering, we designed a three-dimensional evaluation metric that automatically assesses questions from the perspectives of fluency, accuracy and evidence. Experimental results show that our designed metrics align strongly with the annotations of operations experts.
The contributions of our paper are as follows:
  • We propose OpsEval, encompassing 8920 questions and designing a semi-automatic annotation method, conducting independent and robust evaluations across eight domains, providing guidance for model selection in the Ops domain.
  • Based on OpsEval, we systematically tested over 17 mainstream LLMs, and the results show that the current models do not perform ideally on some Ops tasks, indicating a significant gap towards real-world applications.
  • We designed a multi-dimensional evaluation metric based on GPT-4 (OpenAI, 2023a), which has shown consistency with expert annotations reaching 92%, replacing BLUE for automated domain-specific QA evaluations.
  • We released an online leaderboard that updates the performance of newly emerging models on Ops tasks. Currently, the size of evaluation dataset on this platform continues to grow.
  • To assist researchers in preliminary evaluating their OpsLLMs, we have carefully selected and released 20% of QAs from our benchmark, with the remaining 80% of undisclosed data preventing unfair evaluations due to data leakage (Wei et al., 2023b). We also provide an interface on our website where researchers can upload their models for evaluation against the full dataset.
2. Related Works
@@
As LLMs evolve rapidly, their complex and varied capabilities are increasingly recognized, while traditional NLP metrics fall short of accurately evaluating these abilities. As a result, there is a growing trend towards proposing evaluation benchmarks tailored specifically for LLMs. These can be divided into two categories: general ability benchmarks and domain-specific benchmarks.
@@
General ability benchmarks assess the general abilities of LLMs across various tasks. These tasks evaluate LLMs’ capacity for logical reasoning, general knowledge, common sense, and other similar abilities rather than being confined to a particular domain. HELM (Liang et al., 2022) employs seven distinct metrics in 42 unique scenarios, offering a comprehensive evaluation of LLMs’ capabilities across multiple dimensions. BIG-bench (Srivastava et al., 2022) comprises 204 tasks spanning a wide array of topics, with a particular focus on tasks deemed beyond the reach of current LLMs. C-Eval (Huang et al., 2023) is the first comprehensive Chinese evaluation suite designed to assess Chinese LLMs’ advanced knowledge and reasoning abilities rigorously. AGIEval (Zhong et al., 2023) curates authentic questions from examinations such as the Chinese College Entrance Exam (CCEE) and the SAT, constructing a fundamentally human-centric evaluation dataset. In a parallel endeavor, MMCU (Zeng, 2023) leverages questions from the CCEE and various professional exams to establish a robust comprehension benchmark. CG-Eval (Zeng et al., 2023) focuses on assessing the generation capabilities of LLMs, employing a testing framework that includes term definitions, short-answering, and computation problems.
Figure 1. The framework of OpsEval
@@
Domain-specific benchmarks evaluate the abilities of LLMs to handle tasks in specific fields. These benchmarks require LLMs to possess specialized knowledge in a specific domain and to respond in a manner consistent with the cognitive patterns of that field. Despite the rapid progression of LLMs in specialized domains, the evaluation metrics for these specific areas have received less attention. FinEval (Zhang et al., 2023) is a benchmark explicitly designed to measure the advanced financial knowledge of Chinese LLMs. MultiMedQA (Singhal et al., 2022) is an extensive medical question-answering dataset, with questions derived from professional medical exams, research, and consultation records. Huatuo-26M (Li et al., 2023) comprises actual medical consultation records and medical knowledge question-answering content. CMB (Wang et al., 2023a) includes multi-choice questions (CMB-Exam) and complex clinical questions based on real case studies (CMB-Clin), with the correct answers established through expert consensus. NetOps (Miao et al., 2023) focuses on evaluations in the network field, which is relevant to the field of Ops. NetOps includes multi-choice questions in both English and Chinese and a few filling-blanks and question-answering questions. However, they only focus on wired network operations. In contrast, our research encompasses various Ops sub-domains. Furthermore, we have delineated the categorization of tasks and abilities, thereby providing a more nuanced and comprehensive assessment framework. As Ops problems in the real world often involve full-stack IT technologies, we believe OpsEval’s broader scope will provide significantly more benefits than NetOps.
3. OpsEval benchmark
@@
Figure 1 shows the overall framework of OpsEval from construction to evaluation. We first collected data from multiple sources, then enhanced and assured the quality of the collected questions through a question curation process. In the third step, we standardized the format of questions that had undergone manual review. Finally, we evaluated the leading large language models on the dataset using various prompt engineering techniques. The following subsections provide detailed descriptions of each step.
3.1. Data Collection
@@
Our benchmark questions have been collected from various sources; we summarize them into four categories: company materials, certification exams, Ops textbooks, and automated generation. Each source is highly esteemed globally and reviewed by our collaborators.
@@
Company Materials include production environment materials like Ops tickets and error logs, as well as internal documents and tests for Ops staff training. We have established cooperative relationships with 10 companies and received materials from them. The companies and enterprises we collaborate with cover various sectors, including Internet, telecommunications, finance, and Ops service/tool providers. Each company has designated experts to work with us to select questions. Information about the companies and experts can be found in the appendix.
@@
Certification Exams include knowledge assessments necessary for becoming an Ops staff and are naturally in the form of multiple-choice and short-answer questions. We obtained the relevant study guidebooks for these certification exams from public book websites and extracted sample questions from them as one of the sources for Ops questions.
@@
Besides the sources described above, to enhance the diversity and depth of our test set, we also source question-answering questions from authoritative books covering a range of Ops domains by extracting textbook contents and providing them to GPT4. In Sec 6, we discussed the methodology and challenges in Automated Generation, hoping to provide some experience in Ops benchmark creation.
By the end of the data collection process, we have collected 21250 questions in total, which are to be further selected and curated.
3.2. Quality Enhancement
We systematically carried out the processing of our original test set in the following stages:
Deduplication: Any repeated or highly similar questions are identified and removed to avoid redundancy in the test set. We calculate the cosine similarity of the question stems to detect duplicate questions and identify pairs of questions with a similarity above a certain threshold (th=0.7). These pairs are later manually reviewed to confirm if they are duplicates.
Dependance Filtering: Since OpsEval primarily focuses on the ability of large language models to understand Ops knowledge, we have filtered out questions that rely on external images or document content to ensure the completeness of the question content itself.
Question Categorization: In the complex landscape of Ops, recognizing the multidimensional nature of tasks is essential. We devise a categorization that captures many tasks that professionals confront in practical applications. The categorization process consists of two steps: automated screening and manual review.
We first use GPT-4 for topic modeling to gain rough insights about the dataset, leveraging its capabilities to determine the relevance of each question to Ops. The topic modeling resulted in more than 20 tasks but had an imbalanced distribution. The prompt we used to let GPT-4 do the categorization is shown in Figure 2(a).
Secondly, during the manual review process (as we will introduce this step later in this section), we involved dozens of experts in manual screening. By merging similar tasks and labeling different ability levels, we categorize the questions into eight tasks and three ability levels, respectively.
  • Task Categorization: Our eight distinct Ops tasks are formulated by industry relevance, task frequency, and the significance of each area in Ops. Details of the eight tasks can be found in Appendix A.2.
  • Ability Categorization: Based on which ability is required to answer them, questions are classified into three categories: Knowledge Recall, Analytical thinking, and Practical Application, reflecting the challenges professionals might encounter in real-world scenarios. The three abilities are described in Appendix A.3.
Table 1. The distribution of different tasks and abilities of questions in OpsEval.

### v4

# OpsEval: A Comprehensive Benchmark Suite for Evaluating Large Language Models’ Capability in IT Operations Domain

@@
Abstract
The past decades have witnessed the rapid development of Information Technology (IT) systems, such as cloud computing, 5G networks, and financial information systems. Ensuring the stability of these IT systems has become an important issue. Large language models (LLMs) that have exhibited remarkable capabilities in NLP-related tasks are showing great potential in AIOps, such as root cause analysis of failures, generation of operations and maintenance scripts, and summarizing of alert information. Unlike knowledge in general corpora, knowledge of Ops varies with the different IT systems, encompassing various private sub-domain knowledge, sensitive to prompt engineering due to various sub-domains, and containing numerous terminologies. As a consequence, existing NLP-related benchmarks such as C-Eval and MMLU can not guide the selection of suitable LLMs for Ops, and current metrics like BLEU and ROUGE can not adequately reflect the question-answering (QA) effectiveness in the Ops domain. Therefore, this paper proposes a comprehensive benchmark suite named OpsEval, including an Ops-oriented evaluation dataset, an evaluation benchmark for Ops, and a specially designed QA evaluation method for Ops. Our dataset contains 7,184 multiple-choice questions and 1,736 QA questions. We have carefully selected and released 20% of the dataset by CC-BY-NC-4.0 license [^1] Data page is available at https://github.com/NetManAIOps/OpsEval-Datasets. written by domain experts in various sub-domains to assist current researchers in preliminary evaluations of their LLMs tailored for Ops (OpsLLM). The remaining undisclosed 80% of the data is used to prevent test set leakage. We test over 21 latest LLMs under various settings such as self-consistency, chain-of-thought, and in-context learning, revealing findings when applying LLMs to Ops. We also propose an evaluation method for QA in Ops, which has a coefficient of 0.9175 with human experts and is improved by 0.2470 and 1.313 compared to BLEU and ROUGE, respectively. Over the past seven months, our dataset and leaderboard [^2] Leaderboard is available at https://opseval.cstcloud.cn/content/leaderboard. have been continuously updated. For reproducibility, the evaluation framework code is publicly available.
1 Introduction
@@
In recent years, large language models (LLMs) have witnessed significant advancements. The latest models, such as GPT-4o (OpenAI, 2024), GPT-4V (OpenAI, 2023b), Meta-Llama-3 (AI@Meta, 2024), and GLM-4 (Zeng et al., 2022), have demonstrated exceptional generalization and task-planning capabilities. As a result, these models have provided numerous opportunities to enhance downstream domain-specific applications. With its advanced summarizing, report analyzing, and ability to diagnose errors, LLM is well suited for Ops on tasks like question answering, information summarizing, and report analysis. Hereinafter, we refer to the LLM used for Ops as OpsLLM, regardless of whether they have been optimized specifically for Ops.
@@
While there are benchmarks for assessing general-purpose NLP-related capabilities, such as C-EVAL (Huang et al., 2023) and AGIEval (Zhong et al., 2023), there are also benchmarks for specific domains, like FinEval (Zhang et al., 2023) in the financial sector and CMB (Wang et al., 2023a) in the medical sector. However, no benchmark exists to evaluate the effectiveness of LLMs or OpsLLMs in Ops tasks. There is an urgent need for an Ops benchmark that informs us about the performance of current LLMs on Ops tasks. On the other hand, a good benchmark can significantly aid the optimization process of OpsLLMs tailored for the Ops domain. Nevertheless, due to the specialty of the Ops tasks, constructing an Ops benchmark presents the following challenges:
@@
  • Sensitive data. The Ops data is primarily sensitive and proprietary to companies, with very few publicly available data, making it difficult for any company to independently provide sufficient evaluation data to ensure confidence in the test results.
  • Sub-domains. There are many different sub-domains, such as 5G communications, cloud computing, and bank transaction systems. Ops within each sub-domain typically requires a combination of various capabilities, such as network configuration or terminology explanation, which we refer to as “tasks”. Due to the large number of sub-domains and tasks in the Ops field and the lack of an authoritative and systematic taxonomy, classifying a large number of questions becomes a challenge.
@@
  • Prompt sensitivity. Due to the relatively proprietary nature of the Ops, existing LLMs have not undergone specialized supervised fine-tuning (SFT) for instruct following within the Ops field, the evaluation results are more sensitive to prompt engineering. Designing appropriate prompts for robust and accurate evaluation is challenging.
@@
  • QA metric. Existing metrics like BLEU only consider the similarity of model output to the reference in natural language aspects, which does not always reflect true performance in Ops tasks. Some terms and expressions in Ops sub-domains have specific meanings that LLM cannot summarize. Designing an automatic evaluation method that assesses the accuracy of QA of Ops from an accurate semantic level is challenging.
@@
To address these issues, we propose OpsEval, a comprehensive benchmark suite for evaluating LLMs’ capability in the IT operations domain. First, to tackle the challenge of benchmark data mostly being private and not publicly shareable, we initiated a community around AIOps, which has attracted dozens of companies to participate. We have selected 10 representative sub-domains from the community, allowing continuous data contributions from community members. We then aggregate data under the same sub-domain to ensure robustness in evaluation. Additionally, we generated multi-choice (MC) and question-answering (QA) questions as supplements based on publicly available network management books. To address the challenge of classifying the numerous sub-domains and tasks in the Ops field, we employ model-based pre-clustering and manual review to annotate eight tasks and three abilities for independent evaluation. Considering the prompt sensitivity of benchmark results, we systematically test model performance under self-consistency (SC), chain-of-thought (CoT), and few-shot in-context learning (ICL). The prompts used in our evaluation are also disclosed in the paper. Lastly, to address the inaccuracy of existing metrics in Ops QA evaluation, we design FAE-Score, which assesses questions from the perspectives of fluency, accuracy, and evidence. Experimental results show that our designed metrics align strongly with the annotations of human experts.
The contributions of our paper are as follows: 1) We introduce OpsEval, the first bilingual multi-task dataset in the operations domain, covering 8 tasks and 3 abilities with 8920 questions. To assist researchers in preliminary evaluating their OpsLLMs, we have carefully selected and released 20% of QAs from our benchmark licensed under CC-BY-NC-4.0, with the remaining 80% of undisclosed data preventing unfair evaluations due to data leakage (Wei, et.al., 2023) 2) Based on the dataset, we introduce the OpsEval evaluation benchmark, conducting independent and robust evaluations with various prompting techniques and a specifically designed evaluation metric named FAE-Score. Compared to the commonly employed BLEU and ROUGE metrics, FAE-Score exhibits a more pronounced congruence with the evaluations of human experts. Specifically, FAE-Score attains a correlation coefficient 0.9175 with expert assessments, surpassing the coefficients of 0.6705 for BLEU and -0.3957 for ROUGE. 3) We released an online leaderboard that continuously updates mainstream LLMs’ performance on Ops tasks for the past 7 months. To ensure the reproducibility and reliability of this leaderboard, we made our evaluation framework’s code publicly available [^4] Evaluation framework code is available at https://github.com/NetManAIOps/OpenCompass-OpsQA.. 4) Based on the results of OpsEval evaluation, we provide key observations and practical lessons to help domain practitioners make decisions such as whether existing models are sufficiently applicable within a specific sub-domain, the necessity for fine-tuning and whether model quantization compromises the effectiveness.
2 Related Works
@@
As LLMs evolve rapidly, their complex and varied capabilities are increasingly recognized. As a result, there is a growing trend towards evaluation benchmarks tailored specifically for LLMs. These can be divided into two categories: general ability benchmarks and domain-specific benchmarks.
Table 1: A comparison of OpsEval with other popular datasets/benchmarks.
MMLU HELM BIG-bench SEAL C-Eval AGIEval FLUE MultiMedQA CMB NetOps OpsEval
Ops Domain Dataset
Open-sourced Benchmark
Up-to-date Leaderboard
@@
General ability benchmarks assess the general abilities of LLMs across various tasks. These tasks evaluate LLMs’ capacity for logical reasoning, general knowledge, common sense, and other similar abilities rather than being confined to a particular domain. MMLU (Hendrycks et al., 2021) is a benchmark designed to measure knowledge acquired during pretraining by evaluating models exclusively in zero-shot and few-shot settings, covering 57 subjects across STEM. HELM (Liang et al., 2022) employs seven distinct metrics in 42 unique scenarios, offering a comprehensive evaluation of LLMs’ capabilities across multiple dimensions. BIG-bench (Srivastava et al., 2022) comprises 204 tasks spanning a wide array of topics, with a particular focus on tasks deemed beyond the reach of current LLMs. SEAL (AI, 2024b) features private, expert evaluations of leading frontiers models. C-Eval (Huang et al., 2023) is a comprehensive Chinese evaluation suite designed to assess Chinese LLMs’ advanced knowledge and reasoning abilities rigorously. AGIEval (Zhong et al., 2023) curates authentic questions from examinations such as the Chinese College Entrance Exam (CCEE) and the SAT, constructing a fundamentally human-centric evaluation dataset.
@@
Domain-specific benchmarks evaluate the abilities of LLMs to handle tasks in specific fields. These benchmarks require LLMs to possess specialized knowledge in a specific domain and to respond in a manner consistent with the cognitive patterns of that field. Despite the rapid progression of LLMs in specialized domains, the evaluation metrics for these specific areas have received less attention. FLUE (Shah et al., 2022) is an open-source comprehensive suite of benchmarks, including new benchmarks across 5 NLP tasks in financial domain. MultiMedQA (Singhal et al., 2022) is an extensive medical question-answering dataset, with questions derived from professional medical exams, research, and consultation records. CMB (Wang et al., 2023a) includes multi-choice questions (CMB-Exam) and complex clinical questions based on real case studies (CMB-Clin). NetOps (Miao et al., 2023) focuses on evaluations in the network field, which is relevant to the field of Ops. NetOps includes multi-choice questions in both English and Chinese and a few cloze tests and question-answering questions. However, they only focus on wired network operations and while the dataset is released, they lack a benchmark that continuously updates the leaderboard.
3 OpsEval Benchmark
Figure 1: The framework of OpsEval
@@ 3 OpsEval Benchmark
Figure 1 shows the overall framework of OpsEval from construction to evaluation. We collected data from multiple sources and then preprocessed it to enhance its quality. Finally, we evaluated LLMs on the dataset using various prompt engineering techniques.
3.1 Data Collection
@@
Our benchmark questions have been collected from various sources; we summarize them into four categories: company materials, certification exams, Ops textbooks, and automated generation. Each source is highly esteemed globally and reviewed by our Ops collaborators.
@@
Company Materials. include production environment materials like Ops tickets and error logs , as well as internal documents and tests for Ops staff training. We have established cooperative relationships with 10 companies, covering various sectors like telecommunications, finance, and Ops service/tool providers, and received expert collaboration and Ops materials from them. The Appendix A.1 provides information about the companies and experts.
@@
Certification Exams. include knowledge assessments necessary for becoming an Ops staff and are naturally in the form of multiple-choice and question-answering questions. We obtained the relevant study guidebooks for these certification exams from public book websites and extracted sample questions from them as one of the sources for Ops questions.
@@
Automated Generation. To enhance the diversity and depth of our test set, we source QAs from authoritative books covering a range of Ops domains by extracting textbook contents and asking GPT4 to generate questions. In Appendix A.7, we discussed the methodology and challenges.
Table 2: The number of questions in OpsEval, grouped by their sub-domains.
Sub-domain Source License Type Questions
Wired Network Operation Textbooks Public Multi-Choice 3901
5G Communication Certification Exams Public Multi-Choice 2615
Question-Answering 1162
Oracle Database Company Materials CC-BY-NC-4.0 Multi-Choice 497
Log Analysis Company Materials CC-BY-NC-4.0 Question-Answering 420
DevOps Automated Generation CC-BY-NC-4.0 Question-Answering 154
Securities Information System Company Materials CC-BY-NC-4.0 Multi-Choice 91
Hybrid Cloud Company Materials CC-BY-NC-4.0 Multi-Choice 40
Financial IT Company Materials CC-BY-NC-4.0 Multi-Choice 40
3.2 Preprocessing
We systematically carried out the preprocessing of our original data in the following stages:
Deduplication: Any repeated or highly similar questions are identified and removed to avoid redundancy in the test set. We calculate the cosine similarity of the question stems by bge-large-zh-v1.5  (Xiao et al., 2023) to detect duplicate questions and identify pairs of questions with a similarity above a certain threshold (th=0.7).
Dependance Filtering: We have filtered out questions that rely on external images or document content to ensure the completeness of the question content itself. The filtering process was done by two parallel lists of empirical keywords in the question stems and the responses of GPT-3.5-turbo. The keyword list can be found in the Appendix A.2.
Question Categorization: We devise a categorization that captures many tasks that professionals confront in practical applications. The categorization process consists of two steps: automated screening and manual review. We first use GPT-4 for topic modeling to gain rough insights about the dataset and determine the relevance of each question to Ops, which resulted in more than 20 tasks but had an imbalanced distribution [^5] For prompts used during the preprocessing, please see in Appendix.. We then involved dozens of experts during the manual review process to categorize the questions into eight tasks and three abilities. The distribution of the questions across these eight tasks and three ability levels is shown in Table 3, and the details of each task and ability can be found in Appendix A.3.
Table 3: The distribution of different tasks and abilities of questions in OpsEval.